# Discovery of three magnetic He-sdOs with SALT

### v1

@@
Discovery of three magnetic He-sdOs with SALT
@@
Magnetic fields with strengths ranging from 300 to 500 kG have recently been discovered in a group of four extremely similar helium-enriched hot subdwarf (He-sdO) stars. Besides their strong magnetic fields, these He-sdO stars are characterised by common atmospheric parameters, clustering around $T_{\mathrm{eff}}$ = 46 500 K, $\log g/\mathrm{cm}\,\mathrm{s}^{-1}$ close to 6, and intermediate helium abundances. Here we present the discovery of three additional magnetic hot subdwarfs, J123359.44--674929.11, J125611.42--575333.45, and J144405.79--674400.93. These stars are again almost identical in terms of atmospheric parameters but, at $B\approx$ 200 kG, their magnetic fields are somewhat weaker than those previously known. The close similarity of all known He-sdOs implies a finely-tuned origin. We propose the merging of an He white dwarf with a H+He white dwarf. A differential rotation at the merge interface may initiate a toroidal magnetic field that evolves by a magnetic dynamo to produce a poloidal field. This field is either directly visible at the surface or may diffuse towards the surface if initially buried. We further discuss a broad absorption line centred at about 4630 Å that is common to all magnetic He-sdOs. This feature may not be related to the magnetic field but instead to the intermediate helium abundances in these He-sdO stars, allowing the strong He ii 4686 Å line to be perturbed by collisions with hydrogen atoms.
@@ 1 Introduction
Here we present the discovery of three additional magnetic He-sdOs, J123359.44--674929.11, J125611.42--575333.45, and J144405.79--674400.93, hereinafter J1233--6749, J1256--5753, and J1444--6744, respectively [^2] J1233--6749 $\equiv$ Gaia DR3 5856712448201848320, J1256--5753 $\equiv$ Gaia DR3 6060271804721212672, J1444--6744 $\equiv$ Gaia DR3 5848132374839987840. . These stars closely mirror the atmospheric parameters of the known magnetic He-sdOs, but their average field strengths are slightly lower, between 180 and 220 kG. We also provide a list of six additional He-sdO stars that show a broad absorption feature at about 4630 Å – a feature that is also observed in all magnetic He-sdOs (Pelisoli et al., 2022). A potential non-magnetic origin of this spectral feature is briefly discussed.
@@ 2 Spectroscopic observations
To check for radial velocity and magnetic field variability, we obtained multiple SALT/RSS spectra of the new magnetic He-sdOs J1233--6749, J1256--5753 and J1444--6744. For each of these stars, the first (discovery) spectrum was taken in May 2023 and three follow-up spectra were taken at the end of January 2024, spaced by several days. A summary of these observations is given in Table 2.
Figure 1: He i 4471 Å and He ii 4542 Å for magnetic He-sdOs: previously known stars (top) and new ones from SALT/RSS (bottom). Observations are black and best-fit models red. All spectra are at rest wavelength, offset in steps of 0.3 in normalised flux. They were convolved to the RSS resolution (1.05 $\mathrm{\AA}$), except for that of J0415+2538, which has a worse resolution (1.9 $\mathrm{\AA}$).
Figure 2: Like Fig. 1, but for the unidentified feature, He ii 4686 Å and He i 4713 Å. Lorentz profile fits to the unidentified feature at about 4630 Å are shown in blue.
Table 1: Spectroscopic atmospheric parameters for confirmed magnetic He-sdOs (upper), new magnetic stars (centre), and He-sdOs showing the 4630 Å feature (lower). The last column lists the mean signal-to-noise ratio (S/N) of the RSS spectra. For J1233--6749, J1256--5753, J1444--6744, the four available spectra were co-added to obtain average magnetic field strengths while $T_{\mathrm{eff}}$, $\log g$, and $\log n(\mathrm{He})/n(\mathrm{H})$ were derived from the individual spectra. Estimated systematical uncertainties are stated for one sigma confidence.
@@
333 from Pelisoli et al. (2022). from Dorsch et al. (2022). J0714--2245 $\equiv$ Gaia DR3 2927944940172059648, J1346--4025 $\equiv$ UCAC4 248-062759, J1946--4756 $\equiv$ GALEX J194618.8-475617.
@@ 3 Spectral analysis
A direct way to identify a magnetic hot subdwarf is to resolve line splitting by the linear Zeeman effect. Hydrogen and helium lines are split into three components if the field is strong enough and appear broadened if the field is weaker. In addition, all of the previously known iHe-sdOH stars show a strong feature at about 4630 Å. The origin of this remains unidentified (Pelisoli et al., 2022, their fig. 5). We visually inspected the RSS spectra of all stars obtained as part of the SALT survey so far to identify possibly magnetic He-sdOs using both of these features.
We then used the model grid of Dorsch et al. (2022) to perform $\chi^{2}$ fits [^4] These fits were performed in the isis framework (Houck & Denicola, 2000) by the method of Irrgang et al. (2014). to these spectra and thus obtained best-fit magnetic field strengths, as well as atmospheric parameters $T_{\mathrm{eff}}$, $\log g$ and $\log n(\mathrm{He})/n(\mathrm{H})$. The best-fit parameters for all stars are listed in Table 1. These models are based on Tlusty (Hubeny & Lanz, 2017b) model atmospheres in non-local thermodynamic equilibrium. The inclusion of the magnetic field is simplistic: it is assumed to be homogeneous across the stellar surface and linear Zeeman splitting is only applied during the Synspec (Hubeny & Lanz, 2017a) spectral synthesis calculations [^5] The details are described in Appendix B of Dorsch et al. (2022).. While these models are not a realistic approximation of the actual field geometry, they are sufficient to detect Zeeman splitting and to estimate an average field strength.
@@ 3.1 Three new magnetic He-sdOs
Significant Zeeman splitting was detected in several He i lines and He ii 4686 Å for three stars, J1233--6749, J1256--5753 and J1444--6744, which we consider to be confirmed magnetic. Figs. 1 and 2 show the two most important spectral regions for these stars, He i 4471 Å and the range around He ii 4686 Å, respectively. The Zeeman splitting observed is weaker than in the four previously known magnetic He-sdOs reported by Dorsch et al. (2022) and Pelisoli et al. (2022), and is, in fact, close to the detection limit of RSS. However, there is no other effect that could explain the observed line shapes. The only conceivable other option would be rapid rotation at about 180 km s-1. As shown in Fig. 3, magnetic models provide a much superior fit to our RSS spectra when compared to non-magnetic rotating models. This is because rotation, unlike Zeeman splitting, can only increase the line broadening but not the line strengths.
Figure 3: Detailed view of He ii 4686 Å and He i 4713 Å for the new magnetic He-sdOs. The merged RSS spectra (black) are best matched by Zeeman-split models (red). Models at the best-fit atmospheric parameters but zero magnetic field (dashed red) are excluded even if $v\sin i$ is allowed to vary (blue). At zero field strength, the best-fit $v\sin i$ would equal 190, 180 and 130 km s-1 for J1233--6749, J1256--5753 and J1444--6744, respectively.
Like Pelisoli et al. (2022), we performed $\chi^{2}$ fits using two homogeneous magnetic field components that were allowed to vary in their field strength and contribution, expressed as a surface ratio. These $\chi^{2}$ fits resulted in the lowest average field strengths $B_{\mathrm{av}}$ in hot subdwarfs to date, $222\pm 10$ kG (J1233--6749), $216\pm 24$ kG (J1256--5753) and $184\pm 14$ kG (J1444--6744). These average field strengths were computed by taking the weighted mean of both components $B_{\mathrm{mean}}$ averaged over the four spectra available for each star, as stated in Table 2. Because our spectra were taken with one- and two-day gaps, variation in the magnetic field strength over these four spectra may result from a variable field geometry, possibly caused by rotation combined with magnetic spots. Individual fits are shown in Figs. 9 to 11 for the most important spectral lines; any variability in the field strengths should be confirmed with higher-quality spectra.
@@
The magnetic fields found here are very likely non-homogeneous and more sophisticated models would be needed to further constrain their geometry. Such models would also benefit from a future determination of metal abundances. The strong N iii 4511, 4515 Å lines shown in Fig. 1 seem to suggest that the three new magnetic He-sdOs are significantly enriched in nitrogen compared to the Sun, especially J1444--6744. This star also seems to be enriched in neon as evidenced by strong Ne ii 4290, 4430 Å and Ne iii 4453, 4459 Å lines. Quantitative abundances will be measured by a more detailed analysis in the future.
@@ 3.2 The broad 4630 Å feature
Indeed, the origin of the 4630 Å feature [^6] It should be noted that the 4630 Å feature is in fact not always centred at a rest wavelength of exactly 4630 Å. For the iHe-sdOs considered here, its centre ranges between 4629 Å and 4635 Å at Lorentzian full width at half maxima between 20 Å and 38 Å. may be completely unrelated to the magnetic field. As already discussed by Pelisoli et al. (2022), this feature does not seem to be caused by Zeeman-split metal lines, diffuse interstellar bands, or even ultra-high excitation lines, a feature that is observed for some DO-type white dwarfs (Reindl et al., 2019). Instead, the feature 4630 Å feature seems to be related to the intermediate helium abundances: it could result from collisions with free hydrogen during the formation of He ii 4686 Å. This is similar to the blue satellite lines predicted for the $\mathrm{H}\alpha$ and $\mathrm{H}\beta$ lines in white dwarf stars by Allard et al. (2022) and Spiegelman et al. (2022). However, these features are not detected here, possibly because of the much higher $T_{\mathrm{eff}}$ and lower photospheric densities of intermediate He-sdO stars when compared to the white dwarfs considered by these authors.
@@ 4 Stellar parameters
Figure 5: Hertzsprung-Russell diagram showing known (dark blue) and new magnetic He-sdOs (light blue), as well as 4630Å-stars (violet). Merge tracks from Yu et al. (2021) for a metallicity of $Z=0.01$ and masses of 0.80, 0.65, 0.50 ${\rm M}_{\odot}$ are shown as black lines (solid: core helium burning, dashed: helium shell burning). The pre-helium main sequence phase is shown as a dotted line for 0.65 ${\rm M}_{\odot}$. The helium zero-age main sequence of Paczyński (1971) is shown in red and the grey region represents the EHB. Orange circles represent other He-sdOs from the SALT sample.
@@
The spectral analysis performed in Sect. 3 provides atmospheric parameters, in particular the effective temperature and surface gravity. These can be combined with the Gaia DR3 (Gaia Collaboration et al., 2023) parallax $\varpi$ and an angular diameter $\Theta$ from photometric flux measurements to derive radii $R=\Theta/(2\varpi)$ and luminosities $L=4\pi R^{2}\sigma_{\mathrm{SB}}T^{4}_{\mathrm{eff}}$, where $\sigma_{\mathrm{SB}}$ is the Stefan–Boltzmann constant. To this end we performed spectral energy distribution (SED) fits to all He-sdOs discussed here. As usual, we used Gaia DR3 parallaxes with zero-point corrections applied according to Lindegren et al. (2021) and inflated uncertainties following El-Badry et al. (2021). We could also estimate stellar masses via $M=gR^{2}/G$. These remain poorly constrained due to the high uncertainty in the surface gravity $g$, resulting from additional line broadening introduced by non-homogeneous magnetic fields. To account for such systematic effects, we added estimated uncertainties of 1000 K in $T_{\mathrm{eff}}$, 0.15 in $\log g$ and 0.1 in $\log n(\mathrm{He})/n(\mathrm{H})$ in quadrature to the statistical uncertainties for the computation of stellar parameters. The resulting stellar parameters are listed in Table 3.
@@
While the stellar masses are not well determined, the derived radii and luminosities are quite precise. All magnetic He-sdO stars are located close the zero-age helium main sequence (HeZAMS) in the Hertzsprung-Russell diagram (Fig. 5). This is not uncommon for He-sdO stars, as the comparison to the remaining SALT sample of He-sdO stars shows [^7] A detailed analysis of the full SALT sample will be published in a separate paper.. Indeed, the most likely formation scenario for all (non-binary) He-sdOs is the merging of two white dwarfs, including at least one helium-core WD (Han et al., 2002). The clustering of the now seven confirmed magnetic He-sdO stars at about 46 500 K on the HeZAMS further strengthens the proposition that these stars have formed through a very specific channel. If their masses are interpolated from their location on the HeZAMS of Paczyński (1971) in the Hertzsprung-Russell diagram, their mass distribution is centred at $M^{\mathrm{HeZAMS}}=0.85\pm 0.05$${\rm M}_{\odot}$, which is at the upper end of what is achievable by double HeWD merges (Zhang & Jeffery, 2012). This mass is a rough estimate given that (1) all magnetic He-sdOs retain significant amounts of hydrogen in their atmospheres while the HeZAMS assumes a pure helium composition and (2) the stars may also be somewhat evolved. The interpolated HeZAMS masses for all stars are also listed in Table 3.

### v2

# Discovery of three magnetic helium-rich hot subdwarfs with SALT

@@
Discovery of three magnetic helium-rich hot subdwarfs with SALT
@@
Magnetic fields with strengths ranging from 300 to 500 kG have recently been discovered in a group of four extremely similar helium-rich hot subdwarf (He-sdO) stars. In addition to their strong magnetic fields, these He-sdO stars are characterised by common atmospheric parameters, clustering around $T_{\mathrm{eff}}$ = 46 500 K, a $\log g/\mathrm{cm}\,\mathrm{s}^{-1}$ close to 6, and intermediate helium abundances. Here we present the discovery of three additional magnetic hot subdwarfs, J123359.44--674929.11, J125611.42--575333.45, and J144405.79--674400.93. These stars are again almost identical in terms of atmospheric parameters, but, at $B\approx$ 200 kG, their magnetic fields are somewhat weaker than those previously known. The close similarity of all known He-sdOs implies a finely tuned formation channel. We propose the merging of a He white dwarf with a H+He white dwarf. A differential rotation at the merger interface may initiate a toroidal magnetic field that evolves via a magnetic dynamo to produce a poloidal field. This field is either directly visible at the surface or might diffuse towards the surface if initially buried. We further discuss a broad absorption line centred at about 4630 Å that is common to all magnetic He-sdOs. This feature may not be related to the magnetic field but instead to the intermediate helium abundances in these He-sdO stars, allowing the strong He ii 4686 Å line to be perturbed by collisions with hydrogen atoms.
@@ 1 Introduction
Here we present the discovery of three additional magnetic He-sdOs, J123359.44--674929.11, J125611.42--575333.45, and J144405.79--674400.93, hereinafter J1233--6749, J1256--5753, and J1444--6744, respectively [^2] J1233--6749 $\equiv$ Gaia DR3 5856712448201848320, J1256--5753 $\equiv$ Gaia DR3 6060271804721212672, and J1444--6744 $\equiv$ Gaia DR3 5848132374839987840. . These stars closely mirror the atmospheric parameters of the known magnetic He-sdOs, but their average field strengths are slightly lower, between 180 and 220 kG. We also provide a list of six additional He-sdO stars that show a broad absorption feature at about 4630 Å – a feature that is also observed in all magnetic He-sdOs (Pelisoli et al., 2022). A potential non-magnetic origin of this spectral feature is briefly discussed.
@@ 2 Spectroscopic observations
To check for radial velocity and magnetic field variability, we obtained multiple SALT/RSS spectra of the new magnetic He-sdOs, J1233--6749, J1256--5753, and J1444--6744. For each of these stars, the first (discovery) spectrum was taken in May 2023, and three follow-up spectra were taken at the end of January 2024, spaced by several days. A summary of these observations is given in Table 2.
Figure 1: He i 4471 Å and He ii 4542 Å for magnetic He-sdOs: previously known stars (top) and new ones from SALT/RSS (bottom). Observations are black and best-fit models red. All spectra are at the rest wavelength, offset in steps of 0.3 in normalised flux. They were convolved to the RSS resolution (1.05 $\mathrm{\AA}$) except for that of J0415+2538, which has a worse resolution (1.9 $\mathrm{\AA}$).
Figure 2: Like Fig. 1, but for the unidentified feature, He ii 4686 Å, and He i 4713 Å. Lorentz profile fits to the unidentified feature at about 4630 Å are shown in blue.
Table 1: Spectroscopic atmospheric parameters for confirmed magnetic He-sdOs (upper), new magnetic stars (centre), and He-sdOs showing the 4630 Å feature (lower).
@@
333 from Pelisoli et al. (2022). from Dorsch et al. (2022). J0714--2245 $\equiv$ Gaia DR3 2927944940172059648, J1346--4025 $\equiv$ UCAC4 248-062759, and J1946--4756 $\equiv$ GALEX J194618.8-475617. The last column lists the mean signal-to-noise ratio (S/N) of the RSS spectra. For J1233--6749, J1256--5753, and J1444--6744, the four available spectra were co-added to obtain average magnetic field strengths, while $T_{\mathrm{eff}}$, $\log g$, and $\log n(\mathrm{He})/n(\mathrm{H})$ were derived from the individual spectra. Estimated systematical uncertainties are stated for one sigma confidence.
@@ 3 Spectral analysis
A direct way to identify a magnetic hot subdwarf is to resolve line splitting caused by the linear Zeeman effect. Hydrogen and helium lines are split into three components if the field is strong enough and appear broadened if the field is weaker. In addition, all of the previously known iHe-sdOH stars show a strong feature at about 4630 Å. The origin of this remains unidentified (Pelisoli et al., 2022, their Fig. 5). We visually inspected the RSS spectra of all stars obtained as part of the SALT survey so far to identify possibly magnetic He-sdOs using both of these features.
We then used the model grid of Dorsch et al. (2022) to perform $\chi^{2}$ fits [^4] These fits were performed in the isis framework (Houck & Denicola, 2000) using the Irrgang et al. (2014) method. to these spectra and thus obtain best-fit magnetic field strengths, as well as the atmospheric parameters $T_{\mathrm{eff}}$, $\log g$, and $\log n(\mathrm{He})/n(\mathrm{H})$. The best-fit parameters for all stars are listed in Table 1. These models are based on Tlusty (Hubeny & Lanz, 2017b) model atmospheres in non-local thermodynamic equilibrium. The inclusion of the magnetic field is simplistic: it is assumed to be homogeneous across the stellar surface, and linear Zeeman splitting is only applied during the Synspec (Hubeny & Lanz, 2017a) spectral synthesis calculations [^5] The details are described in Appendix B of Dorsch et al. (2022).. While these models are not a realistic approximation of the actual field geometry, they are sufficient to detect Zeeman splitting and to estimate an average field strength.
@@ 3.1 Three new magnetic He-sdOs
Significant Zeeman splitting was detected in several He i lines and He ii 4686 Å for three stars, J1233--6749, J1256--5753, and J1444--6744, which we consider to be confirmed as magnetic. Figures 1 and 2 show the two most important spectral regions for these stars, He i 4471 Å and the range around He ii 4686 Å, respectively. The Zeeman splitting observed is weaker than in the four previously known magnetic He-sdOs reported by Dorsch et al. (2022) and Pelisoli et al. (2022) and is, in fact, close to the detection limit of RSS. However, there is no other effect that could explain the observed line shapes. The only other conceivable option would be rapid rotation at about 180 km s-1. As shown in Fig. 3, magnetic models provide a much superior fit to our RSS spectra when compared to non-magnetic rotating models. This is because rotation, unlike Zeeman splitting, can only increase the line broadening, not the line strengths.
Figure 3: Detailed view of He ii 4686 Å and He i 4713 Å for the new magnetic He-sdOs. The merged RSS spectra (black) are best matched by Zeeman-split models (red). Models with the best-fit atmospheric parameters but zero magnetic field (dashed red) are excluded even if $v\sin i$ is allowed to vary (blue). At zero field strength, the best-fit $v\sin i$ would equal 190, 180, and 130 km s-1 for J1233--6749, J1256--5753, and J1444--6744, respectively.
Like Pelisoli et al. (2022), we performed $\chi^{2}$ fits using two homogeneous magnetic field components that were allowed to vary in terms of field strength and contribution, expressed as a surface ratio. These $\chi^{2}$ fits resulted in the lowest average field strengths, $B_{\mathrm{av}}$, in hot subdwarfs to date, $222\pm 10$ kG (J1233--6749), $216\pm 24$ kG (J1256--5753), and $184\pm 14$ kG (J1444--6744). These average field strengths were computed by taking the weighted mean of both components $B_{\mathrm{mean}}$ averaged over the four spectra available for each star, as stated in Table 2. Because our spectra were taken with one- and two-day gaps, the variation in the magnetic field strength between these four spectra may result from a variable field geometry, possibly caused by rotation combined with magnetic spots. Individual fits are shown in Figs. 9 to 11 for the most important spectral lines; any variability in the field strengths should be confirmed with higher-quality spectra.
@@
The magnetic fields found here are very likely non-homogeneous, and more sophisticated models would be needed to further constrain their geometry. Such models would also benefit from a future determination of metal abundances. The strong N iii 4511, 4515 Å lines shown in Fig. 1 seem to suggest that the three new magnetic He-sdOs are significantly enriched in nitrogen compared to the Sun, especially J1444--6744. This star also seems to be enriched in neon, as evidenced by strong Ne ii 4290, 4430 Å and Ne iii 4453, 4459 Å lines. Quantitative abundances will be measured via a more detailed analysis in the future.
@@ 3.2 The broad 4630 Å feature
The origin of the 4630 Å feature [^6] It should be noted that the 4630 Å feature is in fact not always centred at a rest wavelength of exactly 4630 Å. For the iHe-sdOs considered here, its centre ranges between 4629 Å and 4635 Å at Lorentzian full widths at half maximum between 20 Å and 38 Å. may be completely unrelated to the magnetic field. As already discussed by Pelisoli et al. (2022), this feature does not seem to be caused by Zeeman-split metal lines, diffuse interstellar bands, or even ultra-high excitation lines, a feature that is observed for some DO-type WDs (Reindl et al., 2019). Instead, the 4630 Å feature seems to be related to the intermediate helium abundances: it could result from collisions with free hydrogen during the formation of He ii 4686 Å. This is similar to the blue satellite lines predicted for the $\mathrm{H}\alpha$ and $\mathrm{H}\beta$ lines in WD stars by Allard et al. (2022) and Spiegelman et al. (2022). However, these features are not detected here, possibly because of the much higher $T_{\mathrm{eff}}$ and lower photospheric densities of intermediate He-sdO stars when compared to the WDs considered by Allard et al. (2022) and Spiegelman et al. (2022).
@@ 4 Stellar parameters
Figure 5: Hertzsprung-Russell diagram showing known (dark blue) and new magnetic He-sdOs (light blue), as well as 4630Å  stars (violet). Merger tracks from Yu et al. (2021) for a metallicity of $Z=0.01$ and masses of 0.80, 0.65, and 0.50 ${\rm M}_{\odot}$ are shown as black lines (solid for core helium-burning and dashed for helium shell-burning). The pre-helium main sequence phase is shown as a dotted line for 0.65 ${\rm M}_{\odot}$. The Paczyński (1971) HeZAMS is shown in red, and the grey region represents the extreme horizontal branch. Orange circles represent other He-sdOs from the SALT sample.
@@
The spectral analysis performed in Sect. 3 provides atmospheric parameters, in particular the effective temperature and surface gravity. They can be combined with the Gaia Data Release 3 (DR3; Gaia Collaboration et al., 2023) parallax, $\varpi$, and an angular diameter, $\Theta,$ from photometric flux measurements to derive radii $R=\Theta/(2\varpi)$ and luminosities $L=4\pi R^{2}\sigma_{\mathrm{SB}}T^{4}_{\mathrm{eff}}$, where $\sigma_{\mathrm{SB}}$ is the Stefan–Boltzmann constant. To this end, we performed spectral energy distribution (SED) fits to all the He-sdOs discussed here. As usual, we used Gaia DR3 parallaxes with zero-point corrections applied according to Lindegren et al. (2021) and inflated uncertainties following El-Badry et al. (2021). We could also estimate stellar masses via $M=gR^{2}/G$. These masses remain poorly constrained due to the high uncertainty in the surface gravity, $g$, resulting from additional line broadening introduced by non-homogeneous magnetic fields. To account for such systematic effects, we added estimated uncertainties of 1000 K in $T_{\mathrm{eff}}$, 0.15 in $\log g$, and 0.1 in $\log n(\mathrm{He})/n(\mathrm{H})$ in quadrature to the statistical uncertainties for the computation of stellar parameters. The resulting stellar parameters are listed in Table 3.
@@
While the stellar masses are not well determined, the derived radii and luminosities are quite precise. All magnetic He-sdO stars are located close the zero-age helium main sequence (HeZAMS) in the Hertzsprung-Russell diagram (Fig. 5). This is not uncommon for He-sdO stars, as the comparison to the remaining SALT sample of He-sdO stars shows [^7] A detailed analysis of the full SALT sample will be published in a separate paper.. The most likely formation scenario for all (non-binary) He-sdOs is the merging of two WDs, at least one of which is a helium-core WD (Han et al., 2002). The clustering of the now seven confirmed magnetic He-sdO stars at about 46 500 K on the HeZAMS further strengthens the proposition that these stars formed through a very specific channel. If their masses are interpolated from their location on the Paczyński (1971) HeZAMS in the Hertzsprung-Russell diagram, their mass distribution is centred at $M^{\mathrm{HeZAMS}}=0.85\pm 0.05$${\rm M}_{\odot}$, which is at the upper end of what is achievable by double He-WD mergers (Zhang & Jeffery, 2012). This mass is a rough estimate given that (1) all magnetic He-sdOs retain significant amounts of hydrogen in their atmospheres, while the HeZAMS assumes a pure helium composition and (2) the stars may also be somewhat evolved. The interpolated HeZAMS masses for all stars are also listed in Table 3.